A team of three men receive a check for $9200. The Journeymen makes 300% more than what the Apprentice makes and the Master Craftmen makes 250% more of what the Journeymen makes. How much does each man make?
1
Expert's answer
2011-06-30T07:05:45-0400
Let's denote the amount of& money for Apprentice as X,
hence the Journeymen receives (100% + 300%)/100% X = X + 3 X = 4 X,
and the Master Craftmen gets (100% + 250%)/100% * (4X) = 3.5 * 4X = 14X.
Since the team receives $9200, X + 4X + 14X = 19X = $9200
X = 9200/19 = $484.21.
The Apprentice makes $484.21, the Journeymen makes $1936.84, the Master Craftmen -& $6778.95.